# Equilibrium-distribution-function based mesoscopic finite-difference methods for partial differential equations: Modeling and Analysis

### v1

@@ III.5 The equilibrium, and source distribution functions of MesoDF method for NCDE and NSEs
Remark 2 : For incompressible flow, inserting $\rho=\rho_{0}+\delta\rho$ with $\delta\rho=O(Ma^{2})$ into Eq. (111), and omitting the terms of $O(Ma^{3})$, one can obtain the He-Luo-like version [42] of the present MesoFD scheme with $f_{j}^{eq}=g_{j}^{eq}(\rho,\rho_{0}\mathbf{u},\rho_{0}\mathbf{uu})$.
@@ IV Numerical tests
In this section, several different numerical examples are used to test the present MesoFD scheme. Here we only use the MMFD scheme (49). To test the accuracy of the MesoFD scheme, the following global relative errors (GRE) are adopted:
@@
$\phi(x,y,t)=\exp\left[\left(1-2\pi^{2}\kappa\right)t\right]\sin(\pi(x+y)),$
and we also have the initial condition
$\phi(x,y,0)=\sin(\pi(x+y)).$
The computation domain is $(x,y)\in[0,2]\times[0,2]$ and $t\in[0,T]$.
@@
In the simulation, we use the D2Q9 lattice model with $\omega_{0}=4/9,~{}\omega_{1-4}=1/9,~{}\omega_{5-8}=1/36$, and $c_{s}^{2}=c^{2}/3,~{}c=\Delta x/\Delta t$. Then we consider the following cases:
@@
Case 2: Let $A_{10}=A_{11}=-B_{00}=-1,~{}A_{21}=0,~{}A_{22}=\frac{1}{2},~{}a_{0}=\frac{1}{4}$, then we have
@@
Let $N_{x}=\frac{1}{\Delta x}$, $N_{t}=\frac{1}{\Delta t}$, then the errors and convergency orders at $T=1.0$ are listed in Table 1. From the table, the second convergency order can be found, which is in agreement with the theoretical analysis.
@@
Example 2: Isotropic convection-diffusion equation without source term
@@
where $u_{x}=u_{y}=1.0$ and $D=0.02$ is the diffusion coefficient. The computation domain is $(x,y)\in[0,2]\times[0,2]$ and $t\in[0,T]$. The exact solution is
$\phi(x,y,t)=\exp\left(-2D\pi^{2}t\right)\cos[\pi(x+y)-\pi(u_{x}+u_{y})t],$
and we also have the initial condition
$\phi(x,y,0)=\cos(\pi(x+y)).$
We take
@@
Case 2: Let $A_{10}=A_{11}=-1,~{}A_{21}=\frac{1}{4},~{}A_{22}=\frac{1}{2},~{}a_{0}=0,~{}\beta=1$, then we have
@@
Case 3: Let $A_{10}=A_{11}=-1,~{}A_{21}=0,~{}A_{22}=1,~{}a_{0}=0,~{}\beta=\frac{1}{2}$, then we have
@@
In the simulations, $N_{x}=\frac{1}{\Delta x}$, $N_{t}=\frac{1}{\Delta t}$, then the errors and convergency orders at $T=1.0$ are listed in Table 2. From the table, the second convergency order also can be found, which is in agreement with the theoretical analysis. In order to test the stable condition, that is,
$\frac{\mathbf{u}^{2}}{c_{s}^{2}}\leq\frac{1}{1-\omega_{0}}=\frac{9}{5}.$
for the two-level scheme. We use case 1 to verify the stable condition, and take $D=0.01,~{}u_{x}=u_{y}=1.0$. The errors are list in Table 3. From the tabel, we can see that if $\frac{\mathbf{u}^{2}}{c_{s}^{2}}$ is more than 9/5, then the scheme is unstable which is agreement with the Theorem 1.
Table 3: The errors with different $c_{s}^{2}$
@@
The initial conditions are
$\phi(x,y,0)=e^{x+y},\quad\frac{\partial\phi(x,y,0)}{\partial t}=e^{x+y}.$
In the simulation, we take $N_{x}=N_{y}$ and the space step is $\Delta x=\Delta y=1/N_{x}$. We consider two lattice models D2Q5 and D2Q9. Based on the idea of the finite difference method, we have the approximation of $\phi(x,\Delta t)$
@@
We take the equilibrium function as follows
$f_{j}^{eq}=\omega_{j}\phi\left[1+\frac{(\beta c_{s}^{2}\mathbf{I}-c_{s}^{2}\mathbf{I}):(\mathbf{c}_{j}\mathbf{c}_{j}-c_{s}^{2}\mathbf{I})}{2c_{s}^{4}}\right].$
@@
Case 1: We use the D2Q5 model and let $\beta=1,\gamma=\frac{3}{8\beta}=\frac{3}{8}$, we have
@@
Case 2: We use the D2Q9 model and the parameters are the same as in Case 1.
Case 3: We use the D2Q9 model and take $\beta=\frac{3}{4},~{}\gamma=\frac{3}{8\beta}=\frac{1}{2}$, we obtain
@@
It is noted that the scheme in Case 1 is the same as the explicit finite-difference scheme. The errors and convergency orders are listed in Table 4. The second-order convergency can be found and we can obtain some different schemes which cannot be derived from finite-difference method directly.
Table 4: The errors and convergency orders for Case 1-3 of the wave equation
@@ V Conclusions
By utilizing the equilibrium distribution function to represent the distribution function and/or the nonequilibrium distribution function, a new kind of the mesoscopic numerical method (MesoNM), namely the mesoscopic finite difference method (MesoFDM) with a multi-level and multi-point structure is presented for PDEs including the NSEs and NCDE. The macroscopic version of the MesoFD scheme, the MMFD scheme for the conserved variables, can be obtained directly from the moments of the MesoFD scheme.
A detailed DTE analysis on the present MesoFDM is conducted, and the macroscopic moment equations with truncation errors of different orders can be obtained. The NSEs and NCDE can be recovered from these moment equations to second-order accuracy in space. The MMFD scheme is a unified FD scheme. The existing FD schemes derived from the LB models and even the traditional FD schemes are the special cases of the present MMFD scheme. The MesoFDM not only establishes a connection between the MesoNM and the traditional macroscopic FDM, but also extends the LB method and traditional FDM. We can combine the ’bottom-up’ mesoscopic method with the ’top-down’ macroscopic method to design more effective numerical methods by using their own advantages.
@@ Acknowledgements
This work was financially supported by the National Natural Science Foundation of China (Grants No. 12072127, No. 51836003 and No. 12202130).

### v2

@@ III.5 The equilibrium, and source distribution functions of MesoDF method for NCDE and NSEs
Remark 4 : For incompressible flow, inserting $\rho=\rho_{0}+\delta\rho$ with $\delta\rho=O(Ma^{2})$ into Eq. (121), and omitting the terms of $O(Ma^{3})$, one can obtain the He-Luo model [42] of the present MesoFD scheme with $f_{j}^{eq}=g_{j}^{eq}(\rho,\rho_{0}\mathbf{u},\rho_{0}\mathbf{uu})$.
@@ IV Numerical tests
In this section, several different numerical examples are used to test the present MesoFD scheme, while for simplicity, we only consider the MMFD scheme (51). To test the accuracy of the MesoFD scheme, the following global relative errors (GRE) are adopted,
@@
$\phi(x,y,t)=\exp\left[\left(1-2\pi^{2}\kappa\right)t\right]\sin(\pi(x+y)).$
The computational domain is $(x,y)\in[0,2]\times[0,2]$ and $t\in[0,T]$.
@@
In the simulations, we adopt the D2Q9 lattice model with $\omega_{0}=4/9,~{}\omega_{1-4}=1/9,~{}\omega_{5-8}=1/36$, and $c_{s}^{2}=c^{2}/3,~{}c=\Delta x/\Delta t$. We consider the following cases:
@@
Case 2: Let $A_{10}=A_{11}=-B_{00}=-1,~{}A_{21}=0,~{}A_{22}=\frac{1}{2},~{}a_{0}=\frac{1}{4}$, and we can derive
@@
Let $N_{x}=\frac{1}{\Delta x}$, $N_{t}=\frac{1}{\Delta t}$, then the errors and convergency orders at $T=1.0$ are listed in Table 1. From the table, the second-order convergency can be found, which is in agreement with the theoretical analysis.
@@
Example 2: Isotropic convection-diffusion equation without the source term
@@
where $u_{x}=u_{y}=1.0$ and $D=0.02$ is the diffusion coefficient. The computational domain is $(x,y)\in[0,2]\times[0,2]$ and $t\in[0,T]$, and the exact solution can be expressed as
$\phi(x,y,t)=\exp\left(-2D\pi^{2}t\right)\cos[\pi(x+y)-\pi(u_{x}+u_{y})t].$
Similar to above discussion, we also take
@@
Case 2: Let $A_{10}=A_{11}=-1,~{}A_{21}=\frac{1}{4},~{}A_{22}=\frac{1}{2},~{}a_{0}=0,~{}\beta=1$, then we obtain
@@
Case 3: Let $A_{10}=A_{11}=-1,~{}A_{21}=0,~{}A_{22}=1,~{}a_{0}=0,~{}\beta=\frac{1}{2}$, then we derive
@@
In our simulations, $N_{x}=\frac{1}{\Delta x}$, $N_{t}=\frac{1}{\Delta t}$, and the errors and convergency orders at $T=1.0$ are measured in Table 2. As shown in this table, all of the three cases are of a second-order convergency also can be found, which agrees well with the theoretical analysis. In order to test the stable condition of the two-level scheme, i.e.,
$\frac{\mathbf{u}^{2}}{c_{s}^{2}}\leq\frac{1}{1-\omega_{0}}=\frac{9}{5},$
the case 1 is adopted with $D=0.01,~{}u_{x}=u_{y}=1.0$. We calculate the errors, and list them in Table 3. From the tabel, we can see that if $\frac{\mathbf{u}^{2}}{c_{s}^{2}}$ is more than 9/5, the scheme is unstable which is in agreement with Theorem 1.
Table 3: The errors under different values of $c_{s}^{2}$
@@
In the following simulations, we take $N_{x}=N_{y}$, and consider D2Q5 and D2Q9 lattice models. Based on the idea of the finite difference method, we have the following approximation to the wave equation,
@@
The equilibrium function is taken as follows
$f_{j}^{eq}=\omega_{j}\phi\left[1+\frac{(\beta c_{s}^{2}\mathbf{I}-c_{s}^{2}\mathbf{I}):(\mathbf{c}_{j}\mathbf{c}_{j}-c_{s}^{2}\mathbf{I})}{2c_{s}^{4}}\right]=\omega_{j}\phi\left[1+\frac{(\beta-1)(|\mathbf{c}_{j}|^{2}-2c_{s}^{2})}{2c_{s}^{2}}\right].$
@@
Case 1: Take the D2Q5 model and let $\beta=1,\gamma=\frac{3}{8\beta}=\frac{3}{8}$, we have
@@
Case 2: Consider the D2Q9 model and the parameters are the same as in Case 1.
Case 3: Consider the D2Q9 model and take $\beta=\frac{3}{4},~{}\gamma=\frac{3}{8\beta}=\frac{1}{2}$, we obtain
@@
It is noted that the scheme in Case 1 is an explicit finite-difference scheme. We measure the errors and convergency orders in Table 4, and from which second-order convergency can be found.
Table 4: The errors and convergency orders of Case 1-3 for the wave equation
@@ V Conclusions
By utilizing the equilibrium distribution function to represent the distribution function and/or the nonequilibrium distribution function, a new kind of the mesoscopic numerical method named the mesoscopic finite difference method, with a multi-level and multi-point structure is presented for some PDEs, including but not limited to the NSEs and NCDE. The macroscopic version of the MesoFD scheme, the MMFD scheme for the conserved variables, can be obtained directly from the moments of the MesoFD scheme.
A detailed DTE analysis on the present MesoFDM is conducted at the acoustic scaling, and the macroscopic moment equations (MEs) with truncation errors of different orders can be obtained. Based on the MEs with second- and third-order truncation errors, the NSEs, NCDE and wave equation can be recovered with the second-order accuracy in space. If the MEs with a higher-order truncation errors are used, one can recover the related PDEs with the higher-order accuracy, and more lattice velocities are needed. The MMFD scheme is a unified FD scheme for PDEs in any dimensions. The existing FD schemes derived from the LB models and even the traditional FD schemes are the special cases of present MMFD scheme. The MesoFDM not only establishes a connection between the MesoNM and the traditional macroscopic FDM, but also extends the LB method and traditional FDM. We can combine the ’bottom-up’ mesoscopic method with the ’top-down’ macroscopic method to design more effective numerical methods by using their own advantages.
@@ Acknowledgements
This work was financially supported by the National Natural Science Foundation of China (Grants No. 12072127 and No. 12202130) and Jiangsu Provincial Scientific Research Center of Applied Mathematics, China (No. BK20233002).